OpenStudy (anonymous):

If you can answer this your freakin smart Which two points lie on a line with an undefined slope? A. (3,0) and (2,0) B. (-2,0) and (0,-2) C. (0,0) and (2,2) D. (-2,0) and (-2,5)

OpenStudy (johnweldon1993):

undefinded slope means a vertical line...meaning 'y' = 0 so which is it?

OpenStudy (tyteen4a03):

The equation of a slope is \(\frac{y2-y1}{x2-x1}\). Anything that makes x2 - x1 zero will make the slope undefined, since you cannot divide by zero.

OpenStudy (anonymous):

is it C :o

OpenStudy (anonymous):

undefined slope means both the first coordinates are the same

OpenStudy (luigi0210):

we all know it's that one letter with the points :P \[\frac{ 5-0 }{ -2+2 }=\frac{ 5 }{ 0 }\]

OpenStudy (johnweldon1993):

Yes thank you @satellite73 bad choice of words

OpenStudy (anonymous):

sooo whos gunna just say and earn this Best response metal??

OpenStudy (anonymous):

*medal

OpenStudy (luigi0210):

I'd prefer metal

OpenStudy (johnweldon1993):

not me...lol give it to @Luigi0210 lol

OpenStudy (tyteen4a03):

@Luigi0210 Oooh, gold please!

OpenStudy (luigi0210):

Make it gold :D

OpenStudy (anonymous):

b?

OpenStudy (johnweldon1993):

Not quite...look at what Luigi has written.....and what Satellite has said...an undefined slope will have the same first coordinates.....So it is...?

OpenStudy (anonymous):

ooooh it D isnt it :D

OpenStudy (johnweldon1993):

yes it is!

OpenStudy (luigi0210):

winner winner chicken dinner!

OpenStudy (anonymous):

Grrreat I couldnt of did it without you guys...Thanks a bunch ^.^